The pKa for acetic acid is 4.76. If the concentration of salt in 2mmol/L and that of acetic acid is 6mmol/L, what is the expected pH?

Answer options: (A) 4.43 ; (B) 6.19 ; (C) 104 ; (D) 56

I keep getting pH=4.23 which is not one of the answer options. Can someone walk me through this Q?
Answer #1

You are very close to the answer, and I believe the options are incorrect. I got to the answer by using the formula below.

Use pH = pKa + log ([salt]/[acetic acid])
pH = 4.76 + log ([2 mmol/L]/[6 mmol/L])
pH = 4.76 + (log 1 mmol/L - log 3 mmol/L)
pH = 4.76 + ( 0 - 0.477)
pH = 4.76 + (-0.477)
pH = 4.283

This is what I got and I am pretty sure about my answer, however it's not one of the answer options.

Thank you so much. Kindly rate my response. ?